OpenStudy (aroub):

Given xy+y^2=x+2, find d^2y/dx^2

OpenStudy (perl):

first implicit derivative: 1 * y + x * y ' + 2*y * y ' = 1

OpenStudy (cwrw238):

this is implicit differentiation treat y as a function of x use the product rule/chain rule for xy and chain rule for y^2:- (x *dy/dx + y*1) + 2y * dy/dx = 1 now make dy/dx the subject perform similar process to find the second derivative

OpenStudy (aroub):

I'm tryinggggg :/ I found dy/dx=(1-y)/(x+2y) right?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

now apply the quotient rule

OpenStudy (aroub):

I first did it in terms of dy/dx & then substituted dy/dx by (1-y)/(x+2y). Now, a quick question. If you have this: (1-y)((x+2)/(x+2y)) What do you do exactly? You distribute the (1-y) to the numerator only?

OpenStudy (gorv):

it will be easy if u dont take the term to denominator

OpenStudy (gorv):

1*y+x*dy/dx+2y*dy/dx=1

OpenStudy (gorv):

now first differentiate this but dont move any term on right

OpenStudy (gorv):

differentiation of dy/dx=d^2/dx^2

OpenStudy (cwrw238):

oh - I see - good way to do it!

OpenStudy (gorv):

thanks bro :)

OpenStudy (aroub):

I don't think I get what you mean :p => dy/dx(x+2y)+y=1 ??

OpenStudy (cwrw238):

now you differentiate that implicitly as gorv said derivative of dy/dx is d^2y/dx^2

OpenStudy (cwrw238):

so you use the product rule to differentiate dy/dx(x + 2y)

OpenStudy (cwrw238):

so we have dy/dx ( 1 + 2 * dy/dx)) + d^2y/dx^2( x + 2y) + dy/dx = 0

OpenStudy (cwrw238):

now make d^2y/ dx^2 the subject the substitute for dy/dx

OpenStudy (aroub):

It sounds easier, but sorry I'm not getting it :/ I did it your way (which is also the way i would've done it) & it's pretty long :p I just got lost in the simplifications

OpenStudy (cwrw238):

yes - its a bit messy my way but theres no reason why it can't be done

OpenStudy (aroub):

Actually, wait

OpenStudy (cwrw238):

to be honest i havent tried it my way - i'll look at it now

OpenStudy (cwrw238):

have you understood gorv's method?

OpenStudy (aroub):

Brb

OpenStudy (cwrw238):

sorry i gotta go - i'll look at again later

OpenStudy (cwrw238):

must go

OpenStudy (aroub):

Ahhhh, yesssssss!! I got it :DD Woahh, much easier! THANK YOUUU

OpenStudy (aroub):

No need to look at it again! Thankssss! Bye :)

OpenStudy (cwrw238):

bye